OpenStudy (anonymous):

solve the equation. check both sides and only write the real solution(s). radical 2x+3=x A. -1 B. 3 C. -1,3

OpenStudy (anonymous):

square both sides, solve the quadratic, make sure you test each zero to make sure they work in the original equation

OpenStudy (anonymous):

\[2x+3=x^2\] \[x^2-2x-3=0\] \[(x-3)(x+1)=0\] \[x=3, -1\] \[\sqrt{2(3)+3}=3?\] True, x=3 is a solution \[\sqrt{2(-1)+3}=-1?\] False, x=-1 is not a solution

OpenStudy (anonymous):

but 3 would be correct I was just workin on it and i got 3